OpenStudy (pphalke):

A pharmacist wishes to mix a solution that is 4% Minoxidil. She has on hand 40 ml of a 3% solution and she wishes to add some 6% solution to obtain the desired 4% solution. How much of the 6% solution should she add? whats an easy formula for this

OpenStudy (janu16):

are there choices?

OpenStudy (pphalke):

no on the test i have to take there isnt choices. its just asking how much ml the pharmacist needs

OpenStudy (janu16):

We'll assume it's weight %, and that you are starting with all of the 40 ml of the 3% solution so your problem then is this (40ml * (3g/100ml) + xml * (6g/100ml) )/ (40 + x ml) * 100 = 4 % (1.2 g + 0.06 x)* 100 / (40 + x) = 4 120 + 6 x = 160 + 4x 2x = 40 x = 20 ml

OpenStudy (janu16):

i got 20ml

OpenStudy (pphalke):

thanks its one of the problems i get on the test

OpenStudy (janu16):

Np